# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ IV.5 Depth of the resonance band
$\begin{cases}\frac{\ell_{\rm s}}{d}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle-\gamma\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}r\leq R\>,\\[10.0pt] \frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}r=R\>.\end{cases}$ (94)
The first line of Eq. (94) is the Laplace-transformed diffusion equation, and the second line is the boundary condition at $r=R$. Indeed, we assume that the medium is spherical. The first line of Eq. (94) can be written in the more familiar form
@@
where $\beta$ is the Fourier variable for space, analogous to a wavenumber. The solution of Eq. (95) reads
@@
where $j^{(d)}_{\ell}(z)$ is the generalized spherical Bessel function defined in Eq. (116). It mainly depends on the Bessel function $J_{\frac{d-2}{2}}(z)$. Using Eq. (96), the differential problem (94) becomes the following nonlinear system for the pair $(\gamma,\beta)$
@@
The system of equations (97) is graphically solved in Fig. 7(a).
Figure 7: (a) Graphical solution of the diffusion eigenmodes from Eq. (97) in the dual space $(\gamma,\beta)$ for a 2D disk-shaped random point field with $N=500$ scatterers. The parameters are $R=12.6\,\varsigma$ and $\ell_{\rm s}=1.5\,\varsigma$. The black dots are the first four solutions: $n\in\{1,2,3,4\}$. (b) First diffusion eigenmodes $n\in\{1,2,3\}$ in the position space from Eq. (96).
The first equation corresponds to the parabola, and the second equation to the horizontal straight lines. The intersections between both are the eigenmodes of Eq. (94). The corresponding eigenvalues, denoted as $\beta_{n}$ and $\gamma_{n}$, fully characterize the eigenmodes in space and time, respectively. The profiles of the first diffusion eigenmodes in space are shown in Fig. 7(b) based on Eq. (96).
Furthermore, the spectral decomposition of the diffusion problem (94) allows us to write the general solution for the density and its time evolution. Through inverse Laplace transform, one has the expansion
@@
Although the boundary condition in Eq. (97) cannot be solved for $\beta$ in closed form, it is possible to continue the calculation further if the mean free path is sufficiently small ($\ell_{\rm s}\ll R$). Indeed, for a small enough $\ell_{\rm s}$, one may recognize the first order expansion of
$j^{(d)}_{0}\!\left(\beta R+\tfrac{2V_{d-1}}{S_{d}}\beta\ell_{\rm s}\right)=0\>,$ (99)
in the second equation of Eq. (97). Using the approximate effective radius (91), Eq. (99) can be simplified to
$j^{(d)}_{0}(\beta R_{\rm eff})=0\>.$ (100)
Condition (100) obviously expresses the geometrical interpretation of $R_{\rm eff}$ as the point where the extrapolated density vanishes. This interpretation was already discussed near Eq. (88). The roots of Eq. (100) are given by the zeros of $j^{(d)}_{0}(z)$, or, more explicitly, by the zeros of $J_{\frac{d-2}{2}}(z)$. Thus, we obtain the full spectrum of diffusion eigenmodes
@@
According to the first line of Eq. (97), the values of $\gamma$ associated to Eq. (101) read
@@
The result (102) turns out to be an accurate approximation of the sought eigenvalue of Eq. (93). The quantity (102) is also related to the escape rate of the particle from the medium by $\Gamma_{n}=\left|\gamma_{n}\right|v$.
Finally, using the relationship $\gamma=2{k}_{\rm i}$ introduced in Eq. (45), the diffusion rate of Eq. (102) translates into the following estimate
@@
for the position of the resonance band. The result (103) is considerably more accurate than Eq. (68) of Ref. [47] since it accounts for the Robin boundary condition (97) which is more precise than the Dirichlet condition $j^{(d)}_{\ell}(\beta R)=0$. Equation (103) is also more general than Eq. (68) because it predicts other diffusion eigenmodes than the fundamental one.
@@
Figure 8: (a) Cross-sectional view of the resonance density for a 2D disk-shaped random point field with $N=500$ along the vertical axis ${k}_{\rm r}=6\,\varsigma^{-1}$ and averaged over $2^{10}$ configurations. The parameters are $R=12.6\,\varsigma$ and $\ell_{\rm s}=1.5\,\varsigma$. The vertical dotted lines indicate the diffusion eigenmodes given by Eq. (97). (b) Same as panel (a) but in the 3D case with $R=4.92\,\varsigma$ and $\ell_{\rm s}=2.865\,\varsigma$.
@@
The adequacy between the fundamental diffusion mode ($n=1$) and the maximum of the resonance band in Fig. 8 is remarkable. This is especially true in 3D, because the scale parameter $n\sigma R=1.72$ is not much larger than 11 11 compared to what the diffusion approximation discussed in Sec. IV.3 could have required. In fact, the quality of the approximation (103) greatly depends on the expression of $R_{\rm eff}$ and on the boundary condition. Indeed, if the magnitudes of $R$ and $\ell_{\rm s}$ are the same, there is a factor of two between $R$ and $R_{\rm eff}$, and thus a factor of four in $k_{{\rm idiff},n}$, which is significant. This shows that the accuracy of the prediction (103) is not obvious.
Noticeably, the other diffusion eigenmodes do not lead to any perceptible structure in Fig. 8. We could have thought that Eq. (103) predicts the existence of additional bands deeper in the complex plane. However, we could not find any match, not even in much larger point fields, up to $N=5000$. Only the fundamental mode ($n=1$) seems to play a significant role. Furthermore, regarding the 1D case, these higher-order modes do no coincide with the faint bands in Fig. 6(a) of our previous paper [47]. All these observations lead us to conjecture that these higher-order diffusion eigenmodes do not exist in the quantum case.
A likely reason of this absence is that the transport equation (72), on which Eq. (103) is based, only describes the incoherent (phase-independent) contribution to wave transport. Therefore, this equation is unable to predict the full resonance distribution in the complex $k$ plane, because the latter would be intrinsically determined by coherence and constructive interferences in the same way as the energy eigenstates in quantum mechanics. If this hypothesis is confirmed, the effect of coherent contribution would be to smooth out the comb of higher-order diffusion eigenmodes in Fig. 8.
The issue of the resonance distribution could be overcome with advanced methods capable of describing the coherent contribution to wave transport such as the nonlinear sigma model [83]. This field-theoretical method is based on the average of some generating function over the realizations of a random potential. Upon averaging, the linear wave equation in the disordered medium is then cast into a disorder-free nonlinear wave equation which can be solved semiclassically. This method has proven particularly effective in determining the full statistics of certain observables [83, 84, 43].
V Conclusion and perspectives
@@
More specifically, we showed in Sec. III that the peaks in the resonance density at low energy can be predicted by an effective-medium equation for the disorder-averaged wavefunction. This means that, when the particle wavelength is larger than the mean interscatterer distance, the wave only probes the geometric shape of the medium and not the details due to the disorder.
When the wavelength is smaller than the interscatterer distance, the Bethe-Salpeter equation is required to describe the multiple scattering in the medium. In Sec. IV.4, we checked that the average square modulus of the wavefunction in the bulk predicted by the Bethe-Salpeter equation is consistent with numerical simulations. The square modulus of the wavefunction corresponding to a single configuration of scatterers was also considered (see in particular Fig. 1 for $N=10^{4}$ scatterers) in order to determine the extent of fluctuations relative to the average. These simulations revealed the presence of filaments in the square modulus of the wavefunction which we interpreted as a phenomenon similar to quantum scars [61, *Heller2001]. The detailed study of these structures, including their possible links with the multiple-scattering resonances, is deferred to future works.
Finally, in Sec. IV.5, we solved the Bethe-Salpeter equation in the diffusion regime in order to estimate the resonance band. We showed that the fundamental diffusion eigenmode precisely corresponds to the imaginary part of the resonance band, thus confirming that this band is related to the escape time of the particle in the medium. However, the higher-order diffusion modes given by the diffusion equation do not lead to visible structures in the resonance density. A likely reason is that the Bethe-Salpeter equation ignore coherent effects, although they are crucial in determining energy eigenstates. In the future, we plan to use advanced methods accounting for coherent contributions to wave transport, such as the nonlinear sigma model [83, 84, 43], in order to finely characterize the distribution of resonance widths in strongly open systems.
@@ Acknowledgements.
The present results were obtained during D.G.’s doctoral thesis defended in June 2022. The authors are grateful to Pierre Gaspard for useful discussions and for reviewing this manuscript. The authors also thank Arthur Goetschy and Romain Pierrat for their support in this research. D.G. is grateful to Arthur Goetschy for bringing Refs. [83, 84] to his attention. This work was supported by the Belgian National Fund for Scientific Research (F.R.S.-FNRS) as part of the “Research Fellow” (ASP - Aspirant) fellowship program. This work was also supported by the F.R.S.-FNRS as part of the Institut Interuniversitaire des Sciences Nucléaires (IISN) under Grant No. 4.45.10.08. It also received funding from the French National Research Agency (ANR) as part of the “MARS LIGHT” project.
Appendix A Derivation of the disorder-averaged Green function
We present here a derivation of the effective-medium equation that we consider for the disorder-averaged Green function using the diagrammatic method. The matrix $\mathsf{M}(k)^{-1}$ in Eq. (28) can be expanded in Dyson series as
$\hat{Q}(k)=\sum_{p=0}^{\infty}\sum_{i_{1},\ldots,i_{p}}\hat{G}(k)\hat{F}_{i_{p}}\hat{G}(k)\cdots\hat{G}(k)\hat{F}_{i_{2}}\hat{G}(k)\hat{F}_{i_{1}}\hat{G}(k)\>,$ (104)
where $\hat{F}_{i}=F(k)\left|\boldsymbol{\mathrm{x}}_{i}\right\rangle\left\langle\boldsymbol{\mathrm{x}}_{i}\right|$ denotes the projector onto the $i$-th scatterer weighted by the single-scatterer amplitude. The term $p=0$ in Eq. (104) corresponds to the free propagation: $\hat{G}(k)$. Note that two successive indices, for instance $i_{1}$ and $i_{2}$, cannot be equal to each other, since the Green matrix $\mathsf{G}^{+}(k)$ in Eq. (21) is zero along the diagonal. However, terms such as $i_{1}=i_{3}$ or $i_{1}=i_{4}$ are allowed, for instance. An elegant way of visualizing and classifying the different kinds of terms involved in Eq. (104) is to resort to diagrammatic representations. We consider for this purpose the diagrammatic notations given in Fig. 9(a) and inspired by Refs. [27, 52, *Doicu2019a2, *Doicu2019a3].
Figure 9: (a) Diagrams for the free Green function and the scattering vertex. (b) Expansion of the full resolvent $\hat{Q}(k)$ from Eq. (104). The list of diagrams is not exhaustive. The sums implicitly assume that the indices $i_{1},i_{2},\ldots$ are different from each other. Pairs of identical scatterers are highlighted by a black dotted link.
With these notations, some lowest order terms of the series (104) are shown in Fig. 9(b). In our convention, the diagrams should be read from right to left in the chronological order of the scattering events. Of course, this is not essential because of the time reversal symmetry of the Schrödinger equation. The black dotted links in Fig. 9(b) stress the fact that the scatterer encountered by the particle is the same. The existence of links between adjacent scatterers is forbidden by the zero on the diagonal of $\mathsf{G}(k)$. This is because the collision of the particle on a single scatterer is assumed to be exactly described by the scattering amplitude $F(k)$ and does not need any additional correction.
The full zoo of diagrams in Fig. 9(b) is very difficult to treat in general, especially the great herd of linked diagrams. In order to deal with the series (104), we have to neglect those linked diagrams. This approximation is sometimes known as the Twersky approximation [63, 27, 52, *Doicu2019a2, *Doicu2019a3], although it was already present in Foldy’s paper [28]. A similar approximation in the context of continuous disorder is due to Bourret [68, 19, 20, 21]. The Twersky approximation is based on the idea that, for a sufficiently large set of scatterers ($N\rightarrow\infty$), the chance that the particle collides two times the same scatterer is negligible. The only diagrams which are taken into account consider that the particle takes self-avoiding paths. In this regard, the sums over $i_{1},i_{2},i_{3},\ldots$ in Fig. 9(b) implicitly assume that the indices are all different from each other.
One may verify that the number of neglected diagrams is relatively small for $N\rightarrow\infty$. Indeed, at a given order $p$, the total count of diagrams in Fig. 9(b) is $\mathcal{C}_{\rm tot}=N(N-1)^{p-1}$, whereas the count reduces to $\mathcal{C}_{\rm sa}=N!/(N-p)!$ if we limit ourselves to self-avoiding paths under the Twersky approximation [27, 52, *Doicu2019a2, *Doicu2019a3]. The relative difference between these counts behaves as
$\frac{\mathcal{C}_{\rm tot}-\mathcal{C}_{\rm sa}}{\mathcal{C}_{\rm sa}}=\frac{(p-1)(p-2)}{2N}+\operatorname{\mathcal{O}}(p^{4}/N^{2})\>,$ (105)
which obviously tends to zero for $N\rightarrow\infty$. Equation (105) confirms that the diagrams for which the particle visits the same scatterer several times are less numerous than the self-avoiding paths ($\mathcal{C}_{\rm tot}-\mathcal{C}_{\rm sa}\ll\mathcal{C}_{\rm sa}$). This supports the validity of the Twersky approximation under the large $N$ assumption [27, 52, *Doicu2019a2, *Doicu2019a3].
The Twersky approximation considerably simplifies the calculation of the average Green operator $\langle\hat{Q}(k)\rangle$. Indeed, since the positions of the scatterers are supposed to be statistically independent in our model, the configurational averages can be factored as follows
$\begin{split}&\left\langle\hat{G}(k)\sum_{i_{1}}\hat{F}_{i_{1}}\hat{G}(k)\cdots\hat{G}(k)\sum_{i_{p}}\hat{F}_{i_{p}}\hat{G}(k)\right\rangle\\ &=\hat{G}(k)\left\langle\sum_{i_{1}}\hat{F}_{i_{1}}\right\rangle\hat{G}(k)\cdots\hat{G}(k)\left\langle\sum_{i_{p}}\hat{F}_{i_{p}}\right\rangle\hat{G}(k)\\ &=\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\cdots\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\>,\end{split}$ (106)
where one has used several times the result
$\left\langle\sum_{i=1}^{N}\hat{F}_{i}(k)\right\rangle=\left\langle\sum_{i=1}^{N}F(k)\delta(\hat{\boldsymbol{\mathrm{r}}}-\boldsymbol{\mathrm{x}}_{i})\right\rangle=n(\hat{\boldsymbol{\mathrm{r}}})F(k)\>,$ (107)
coming from Eq. (5). Note that $n(\hat{\boldsymbol{\mathrm{r}}})$ does not commute with $\hat{G}(k)$ in general, because the point field has a finite size and the density thus depends on the position $\hat{\boldsymbol{\mathrm{r}}}$. Substituting Eq. (106) into Eq. (104) gives us the average Green operator
$\begin{split}\langle\hat{Q}(k)\rangle&=\hat{G}(k)\\ &+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\\ &+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\\ &+\cdots\>.\end{split}$ (108)
Again, this series can be more easily visualized using diagrams. We introduce to this end the additional notations given in Fig. 10(a) and inspired by Ref. [53].
Figure 10: (a) Diagrams for the effective Green function and the effective vertex. (b) Resummation of the diagrams of the average resolvent, $\langle\hat{Q}(k)\rangle$, under the Twersky approximation. This step is equivalent to Eqs. (108) and (109).
With these notations, the series (108) translates to the first row of Fig. 10(b). In particular, it becomes more obvious that the series can be cast onto the recursive form
$\langle\hat{Q}(k)\rangle=\hat{G}(k)+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\langle\hat{Q}(k)\rangle\>.$ (109)
With a little bit of algebra, Eq. (109) can also be rewritten as
$\langle\hat{Q}(k)\rangle=\frac{1}{\hat{G}(k)^{-1}-n(\hat{\boldsymbol{\mathrm{r}}})F(k)}=\frac{1}{k^{2}-\hat{\boldsymbol{\mathrm{p}}}^{2}-n(\hat{\boldsymbol{\mathrm{r}}})F(k)}\>.$ (110)
Finally, Eq. (110) directly provides the effective-medium equation (39).
Appendix B Scattering by a sphere and resonances

### v3

@@ III.2 Depth of the resonance band
$\begin{cases}\frac{\ell_{\rm s}}{d}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle-\gamma\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}\boldsymbol{\mathrm{r}}\in\mathcal{V}\>,\\[10.0pt] \frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}\boldsymbol{\mathrm{r}}\in\partial\mathcal{V}\>.\end{cases}$ (49)
The first line of Eq. (49) can be written in the more familiar form
@@
where $\beta$ is the Fourier variable for space, analogous to a wavenumber. As in Sec. III.1, we assume that the medium is spherical so that the solution of Eq. (50) reads
@@
where $j^{(d)}_{\ell}(z)$ is the generalized spherical Bessel function defined in Eq. (64). Using Eq. (51), the differential problem (49) becomes the following nonlinear system for the pair $(\gamma,\beta)$
@@
The roots of the system (52), denoted as $\beta_{n}$ and $\gamma_{n}$, fully characterize the diffusion eigenmodes in space and time, respectively. Furthermore, the spectral decomposition of the diffusion problem (49) allows us to write the general solution for the density and its time evolution. Through inverse Laplace transform, one has the expansion
@@
Although the boundary condition in Eq. (52) cannot be solved for $\beta$ in closed form, it is possible to continue the calculation further if the mean free path is sufficiently small ($\ell_{\rm s}\ll R$). Indeed, for a small enough $\ell_{\rm s}$, one may recognize in the second line of Eq. (52) the first-order expansion of
$j^{(d)}_{0}(\beta R_{\rm eff})=0\>,$ (54)
where $R_{\rm eff}$ is the effective radius
$R_{\rm eff}=R+\frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\>.$ (55)
The condition (54) obviously expresses the geometrical interpretation of $R_{\rm eff}$ as the radius where the extrapolated density vanishes. The roots of Eq. (54) are given by the zeros of $j^{(d)}_{0}(z)$, or, more explicitly, by the zeros of $J_{\frac{d-2}{2}}(z)$. Thus, we obtain the full spectrum of diffusion eigenmodes
@@
According to the first line of Eq. (52), the values of $\gamma$ corresponding to Eq. (56) read
@@
The result (57) turns out to be an accurate approximation of the sought eigenvalues of Eq. (44). The quantity (57) is also related to the escape rate of the particle per unit time by $\Gamma_{n}=\left|\gamma_{n}\right|v$ with $n=1$.
Finally, using the relationship (36), the diffusion rate of Eq. (57) translates into the estimate
@@
with $n=1$ for the position of the resonance band. The result (58) is considerably more accurate than our previous estimate, Eq. (68) of Ref. [47], since it accounts for the Robin boundary condition (52) which is more precise than the Dirichlet condition $j^{(d)}_{\ell}(\beta R)=0$. Equation (58) is also more general than our previous estimate because it predicts other diffusion eigenmodes for $n>1$ than the fundamental one.
@@
Figure 4: (a) Cross-sectional view of the resonance density (28) along the vertical axis ${k}_{\rm r}=6\,\varsigma^{-1}$ for a 2D disk-shaped random point field with $N=500$ and using the point scattering model (16). The parameters are thus $R=12.6\,\varsigma$ and $\ell_{\rm s}=1.5\,\varsigma$. The density is averaged over $2^{10}$ disorder configurations. The vertical dotted lines indicate the diffusion eigenmodes given by Eq. (52). (b) Same as panel (a) but in the 3D case with $R=4.92\,\varsigma$ and $\ell_{\rm s}=2.865\,\varsigma$.
@@
The adequacy between the fundamental diffusion mode ($n=1$) and the maximum of the resonance band in Fig. 4 is remarkable. This is especially true in 3D, because the scale parameter $n\sigma R=1.72$ is not much larger than 11 11 compared to what the diffusion approximation could have required. In fact, the quality of the approximation (58) greatly depends on the expression of $R_{\rm eff}$ and on the boundary condition. Indeed, if the magnitudes of $R$ and $\ell_{\rm s}$ were the same, then there would be a factor of two between $R$ and $R_{\rm eff}$, and thus a factor of four in $k_{{\rm idiff},n}$, which is significant. This shows that the accuracy of the prediction (58) is not obvious and relies on the value (55) of the effective radius $R_{\rm eff}$.
Noticeably, the other diffusion eigenmodes do not lead to any perceptible structure in Fig. 4. We could have thought that Eq. (58) predicts the existence of additional bands deeper in the complex plane. However, we did not find any match, not even in much larger point fields, up to $N=5000$. Only the fundamental mode ($n=1$) appears to play a significant role. Furthermore, regarding the 1D case, these higher-order modes do no coincide with the faint bands in Fig. 6(a) of our previous paper [47]. All these observations lead us to conjecture that these higher-order diffusion eigenmodes do not exist in the quantum case. In this regard, we note that similar distributions, which are devoid of the influence of these extra modes, have been observed in the literature [37, 38, 42, 43].
A likely reason of this absence is that the transport equation (39), on which Eq. (58) is based, only describes the incoherent (phase-independent) contribution to wave transport. Therefore, this equation is unable to predict the full resonance distribution in the complex $k$ plane, because the latter would be intrinsically determined by coherence and constructive interferences in the same way as the energy eigenstates in quantum mechanics. Accordingly, the coherent contributions would smooth out the comb of higher-order diffusion eigenmodes in Fig. 4.
In principle, the resonance distribution could be determined with advanced methods capable of describing the coherent contributions to wave transport such as the nonlinear sigma model [74]. This field-theoretical method is based on the average of some generating function over the realizations of a random potential. Upon averaging, the linear wave equation in the disordered medium is then cast into a disorder-free nonlinear wave equation which can be solved semiclassically. This method has proven particularly effective in determining the full statistics of certain observables [74, 75, 43].
IV Conclusion and perspectives
@@
More specifically, we showed in Sec. III.1 that the peaks in the resonance density at low energy ($\left|k\right|\varsigma\lesssim j_{\frac{d-2}{2}}$) are predicted by the effective-medium equation for the disorder-averaged wavefunction. This means that, when the particle wavelength is larger than the mean interscatterer distance, the wave only probes the geometric shape of the medium and not the details due to the disorder.
Finally, in Sec. III.2, we solved the Bethe-Salpeter equation in the diffusion regime to estimate the location of the resonance band at high energy ($\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$). We showed that the fundamental diffusion eigenmode precisely corresponds to the imaginary part of the resonance band, thus confirming that this band is related to the escape time of the particle in the medium. However, the higher-order diffusion modes given by the diffusion equation do not lead to visible structures in the resonance density. A likely reason is that the Bethe-Salpeter equation ignores the coherent effects that should be taken into account to describe the full shape of the resonance density. In the future, we plan to use advanced methods accounting for coherent contributions to wave transport, such as the nonlinear sigma model [74, 75, 43], in order to finely characterize the distribution of resonance widths in strongly open systems.
@@ Acknowledgements.
The present results were obtained during D.G.’s doctoral thesis defended at ULB in June 2022. The authors are grateful to Pierre Gaspard for useful discussions and for reviewing this manuscript. The authors also thank Arthur Goetschy and Romain Pierrat for their support in this research. D.G. is grateful to Arthur Goetschy for bringing Refs. [74, 75] to his attention. This work was supported by the Belgian National Fund for Scientific Research (F.R.S.-FNRS) as part of the “Research Fellow” (ASP - Aspirant) fellowship program. This work was also supported by the F.R.S.-FNRS as part of the Institut Interuniversitaire des Sciences Nucléaires (IISN) under Grant No. 4.45.10.08. It also received funding from the French National Research Agency (ANR) as part of the “MARS LIGHT” project.
Appendix A Scattering by a sphere and resonances